# Einstein connection of nonsymmetric pseudo-Riemannian manifold

### v1

@@ 4 Explicit form of the Einstein connection of almost contact metric manifold
Let $\nabla$ be an Einstein connection satisfying the $f^{2}$-torsion condition (2.27) on an almost contact metric manifold $(M,f,\xi,\eta,g)$, considered as an NGT space $(M,G=g+F,\nabla)$. Then the torsion $T$ is totally skew-symmetric if and only if (4.5) is true, and the following Codazzi-type condition holds:
@@
In this case, the torsion of $\nabla$ is given by
@@
By Theorem 4.2, we have $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$. Let us consider the $f^{2}$-torsion condition under the additional assumption that $T$ is totally skew-symmetric. Applying T(Y,X,Z)=T(Y,Z,X)T(Y,X,Z)=-T(Y,Z,X), from (4.2) we have
@@
Definition 4.5 (see [5]).
An almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$ is said to be almost-nearly cosymplectic if the Levi–Civita covariant derivative of the fundamental 2-form $F(X,Y)=g(X,fY)$ satisfies
$g\bigl((\nabla^{g}_{X}f)Y,Z\bigr)=\tfrac{1}{3}\,dF(fX,fY,Z)+\tfrac{1}{6}\,\eta(Z)\,d\eta(Y,fX)-\tfrac{1}{2}\,\eta(Y)\,d\eta(fZ,X),$ (4.16)
in particular, the vector field $\xi$ is Killing.
@@
Let $(M^{2n+1},f,\xi,\eta,g)$ be an almost contact metric manifold with fundamental 22–form $F$, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$. Then an Einstein connection $\nabla$ on $M$ has a totally skew-symmetric torsion $T$ if and only if $M$ is almost–nearly cosymplectic, that is, condition (4.16) holds. In this case,
$T(X,Y,Z)$ $=-\tfrac{1}{3}\,dF(X,Y,Z)$
$=-\tfrac{1}{3}\bigl\{\eta(X)d\eta(Y,fZ){+}\eta(Y)d\eta(Z,fX){+}\eta(Z)d\eta(X,fY)\bigr\}+\tfrac{1}{4}N_{f}(fX,fY,fZ).$
The next corollaries of Theorem 4.2 explicitly present the Einstein special connection with the $f^{2}$-torsion condition and the totally skew-symmetric torsion.
@@
$T(X,Y,Z)=\tfrac{1}{2}\big\{(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{fZ}F)(fX,Y)-(\nabla^{g}_{fY}F)(fX,Z)\big\}.$ (4.17)
@@
Let $\nabla$ be the special Einstein connection satisfying the $f^{2}$-torsion condition (2.27) on an almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$, considered as an NGT space $(M,G=g+F,\nabla)$. Then the torsion $T$ is totally skew-symmetric if and only if (4.5) is true, and the torsion of $\nabla$ is given by
@@
Substituting the identity (4.19) into (4.17) yields the first equality of (4.18). Taking $Z\to fZ$ and $Y\to fY$ in (4.19) and using the equalities (4.5), $(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z)$, see (4.3), $\nabla^{g}_{\xi}\,F=\nabla^{g}_{X}F(\xi,\,\cdot)=0$, and in view of (2.5), we get the second equality of (4.18):
@@
Conversely, if (4.18) holds, then $T$ is totally skew-symmetric, since $(\nabla^{g}_{X}F)(Y,Z)$ is skew-symmetric in $(Y,Z)$. The conditions $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$ follow from Theorem 4.2. ∎
5 Explicit form of the Einstein connection of nonsymmetric pseudo-Riemannian space
In this section, we present explicitly an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, in particular, of a weak almost Hermitian manifold $(M^{2n},f,g)$. Recall that a Riemannian manifold $(M^{2n},g)$ endowed with a skew-symmetric $(1,1)$-tensor $f$ of maximal rank $2n$ is called a weak almost Hermitian manifold if $f^{2}$ is not necessarily equal to $\mathrm{Id}$. In particular, we define
$f^{2}:=-\mathrm{Id}-\widetilde{Q}$,
where $\widetilde{Q}$ is a self-adjoint $(1,1)$-tensor. If $\nabla^{g}J=0$, then such $(M^{2n},J,g)$ is called a weak Kähler manifold. The $f^{2}$-torsion condition (2.27) in terms of $\widetilde{Q}$ has the following view:
@@ 5 Explicit form of the Einstein connection of nonsymmetric pseudo-Riemannian space
The following lemma is used at the final steps in the proof of Theorem 5.3.
@@
Let $(M^{2n},f,g)$ be a weak almost Hermitian manifold, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,fY)$. If an Einstein connection $\nabla$ on $M$ satisfies the $f^{2}$-torsion condition (2.27), then
$T(fY,fZ,({\rm Id}{-}f^{2})X)=$ $\,T(Y,Z,({\rm Id}{-}f^{2})f^{2}X){-}2(\nabla^{g}_{f^{2}X}\,F)(Y,fZ){+}2(\nabla^{g}_{f^{2}X}\,F)(fY,Z)$
$+dF(Y,Z,\,({\rm Id}-f^{2})f^{2}X)-dF(fY,fZ,\,({\rm Id}-f^{2})X),$ (5.1)
$T(fY,Z,\,({\rm Id}-f^{2})fX)=$ $\,-T(Y,fZ,\,({\rm Id}-f^{2})fX)-T(Y,Z,\,({\rm Id}-f^{4})X)$
$+dF(Y,Z,\,({\rm Id}-f^{4})X)+2\,(\nabla^{g}_{({\rm Id}-f^{2})X}\,F)(Y,Z)$
@@
The following theorem generalizes Theorem 3.1, it is valid for arbitrary $F$ satisfying (2.27), and for a non-degenerate $F$ we can use it to completely determine $T$.
@@
Let $\nabla$ be an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, satisfying the $f^{2}$-torsion condition (2.27), where $g(X,fY)=F(X,Y)$, and let the tensor $P:={\rm Id}-f^{2}$ have rank $2n$. If $fX\neq 0$, then the torsion $T$ of $\nabla$ is given by
@@
For $fX\neq 0$, it is carried out in a similar, but more computationally complex way than the proof of the Theorem 3.1. Applying $Y\to fY$ and $Z\to fZ$ to (2.7), we have
@@
$\qquad-T(\widetilde{Q}Z,X,\widetilde{Q}Y)-T(X,\widetilde{Q}Y,\widetilde{Q}Z)-T(fY,\widetilde{Q}Z,fX)-T(\widetilde{Q}Y,fZ,fX),$ (5.6)
where $\widetilde{Q}:=-f^{2}-{\rm Id}$. From (2.7) with $Y\to-f^{2}Y$, we get
@@
where $\delta_{4}(X,Y,Z)$ is given in Section 6. Using (2.8) and the $f^{2}$-torsion condition, we reduce $\delta_{4}(X,Y,Z)$ to $\delta_{5}(X,Y,Z)$, given in ection 6. Using (5.1) and (5.1), we find
@@
$T(Y,fZ,\widetilde{Q}f^{3}X)$ $+T(fY,Z,\widetilde{Q}f^{3}X)=dF(Y,Z,\widetilde{Q}f^{2}({\rm Id}+f^{2})X)+2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
@@
$+dF(Y,Z,(3\,\widetilde{Q}{+}2\,\widetilde{Q}^{2})f^{2}X){-}dF(Y,Z,\widetilde{Q}f^{2}({\rm Id}+f^{2})X){+}dF(Y,Z,\widetilde{Q}f^{2}X)$
@@
Take almost Hermitian manifolds $(M_{j},J_{j},g_{j})\ (j=1,\dots,m)$, where $J_{j}^{2}=-{\rm Id}_{\,j}$. Consider the Riemannian product $(M,g)=\bigl(M_{1}\times\cdots\times M_{m},\;g_{1}\oplus\cdots\oplus g_{m}\bigr)$, and define the (1,1)(1,1)–tensor $f=\sqrt{\lambda_{1}}\,J_{1}\oplus\cdots\oplus\sqrt{\lambda_{m}}\,J_{m}$. Then $(M,f,g)$ is a weak almost Hermitian manifold with $f^{2}=-\bigoplus_{j}\lambda_{j}\,{\rm Id}_{j}$, and on $TM_{j}$ we have
@@
$\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,\mathrm{Id}_{j},\quad P=\mathrm{Id}-f^{2}=\bigoplus\nolimits_{j}(1+\lambda_{j})\,\mathrm{Id}_{j},\qquad P^{-1}=\bigoplus\nolimits_{j}\frac{1}{1+\lambda_{j}}\,\mathrm{Id}_{j}.$
@@
If an Einstein connection $\nabla$ satisfies the $f^{2}$-torsion condition (2.27), then $T(X,Y)=0$ for $X\in TM_{i},\ Y\in TM_{j},\ i\neq j$. Hence, the torsion tensor splits into the direct sum $T=T|_{TM_{1}}\oplus\cdots\oplus T|_{TM_{m}}$. Substituting $\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,\mathrm{Id}_{j}$ into (5.3), then using
@@
If each $(M_{j},J_{j},g_{j})$ is a Kähler manifold, then $\nabla^{g_{j}}F_{j}=0$ and $dF_{j}=0$. In this case, all torsion components vanish, and the Einstein connection reduces to the Levi-Civita connection. Using (2.5), we can further simplify the last formula as
$T|_{TM_{j}}(Y,Z,X)=\frac{(\lambda-1)(5-3\lambda)}{4\lambda^{2}}\,(\nabla^{g_{j}}_{X}F_{j})(Y,Z)-\Bigl(\frac{3\lambda^{2}-2\lambda+1}{4\lambda^{2}}+\frac{\lambda-1}{4\sqrt{\lambda}}\Bigr)\,(\nabla^{g_{j}}_{Y}F_{j})(Z,X)$
$\qquad-\Bigl(\frac{3\lambda^{2}-2\lambda+1}{4\lambda^{2}}+\frac{\lambda-1}{4\sqrt{\lambda}}\Bigr)\,(\nabla^{g_{j}}_{Z}F_{j})(X,Y)+\Bigl(\frac{1}{\lambda}+\frac{\lambda-1}{2\sqrt{\lambda}}\Bigr)\,(\nabla^{g_{j}}_{J_{j}X}F_{j})(Y,J_{j}Z)$
$\qquad+\Bigl(\frac{\lambda-1}{4\sqrt{\lambda}}{-}\frac{1}{2\lambda}{-}\frac{\lambda-1}{2\lambda^{2}}\Bigr)\,(\nabla^{g_{j}}_{J_{j}Y}F_{j})(J_{j}Z,X)+\Bigl(\frac{\lambda-1}{4\sqrt{\lambda}}-\frac{1}{2\lambda}{-}\frac{\lambda-1}{2\lambda^{2}}\Bigr)\,(\nabla^{g_{j}}_{J_{j}Z}F_{j})(X,J_{j}Y).$
@@
Let $\nabla$ be an Einstein connection of a weak almost Hermitian manifold $(M^{2n},f,g>0)$, considered as a nonsymmetric Riemannian space $(M,G=g+F)$, where $F(X,Y)=g(X,fY)$. If the $f^{2}$-torsion condition (2.27) is valid, then the torsion $T$ of $\nabla$ is given by (5.3).
@@
For a nonsymmetric Riemannian space $(M,G=g+F)$, the tensor $P={\rm Id}-f^{2}$ is positive definite, and therefore non-degenerate. Using (2.5), we can further simplify (5.3) to the form with $\nabla^{g}\,F$-terms.
@@
$T(fX,Y,Z)=-\tfrac{1}{3}\,dF(fX,Y,Z)=-\tfrac{1}{4}\,N_{f}(X,Y,Z),\quad\nabla\,f^{2}=\nabla^{g}f^{2}=0.$

### v2

@@ 3 Einstein connection of an almost contact metric manifold
Under the conditions of Theorem 3.3, the torsion $T$ is totally skew-symmetric if and only if (3.7) is true, and the following Codazzi-type condition holds:
@@
In this case, the torsion is given by
@@
By Theorem 3.3, we have $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$. Let’s consider the $f^{2}$-torsion condition under the additional assumption that $T$ is totally skew-symmetric. Applying T(Y,X,Z)=T(Y,Z,X)T(Y,X,Z)=-T(Y,Z,X), from (3.3) we have
@@
An almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$ is said to be almost-nearly cosymplectic, see [5], if the Levi–Civita covariant derivative of $f$ satisfies
$g\bigl((\nabla^{g}_{X}f)Y,Z\bigr)=\tfrac{1}{3}\,dF(fX,fY,Z)+\tfrac{1}{6}\,\eta(Z)\,d\eta(Y,fX)-\tfrac{1}{2}\,\eta(Y)\,d\eta(fZ,X),$
where $F(X,Y)=g(X,fY)$, in particular, $\xi$ is a Killing vector field.
@@
Let $(M^{2n+1},f,\xi,\eta,g)$ be an almost contact metric manifold with fundamental 22–form $F$, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$. Then an Einstein connection $\nabla$ on $M$ has a totally skew-symmetric torsion $T$ if and only if the manifold is almost–nearly cosymplectic, in this case, $T(X,Y,Z)=-\tfrac{1}{3}\,dF(X,Y,Z)$.
The next consequences of Theorem 3.3 explicitly present the special Einstein connection with the totally skew-symmetric torsion satisfying the $f^{2}$-torsion condition.
@@
$2\,T(X,Y,Z)=(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{fZ}F)(fX,Y)-(\nabla^{g}_{fY}F)(fX,Z).$ (3.24)
@@
Under the conditions of Theorem 3.3, let $\nabla$ be a special Einstein connection. Then the torsion $T$ is totally skew-symmetric if and only if (3.7) is true; in this case,
@@
Substituting the identity (3.26) into (3.24) yields the first equality of (3.25). Taking $Z\to fZ$ and $Y\to fY$ in (3.26) and using the equalities (3.7), $(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z)$, see (3.4), and $\nabla^{g}_{\xi}\,F=\nabla^{g}_{X}F(\xi,\,\cdot)=0$, in view of (2.5), we get the second equality of (3.25):
@@
Conversely, if (3.25) holds, then $T$ is totally skew-symmetric, since $(\nabla^{g}_{X}F)(Y,Z)$ is skew-symmetric in $Y,Z$. The conditions $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$ follow from Theorem 3.3. ∎
4 Einstein connection of an NGT space
In this section, we present explicitly an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, in particular, of a weak almost Hermitian manifold.
A pseudo-Riemannian manifold $(M^{2n},g)$ endowed with a skew-symmetric (1,1)-tensor $f$ of maximal rank $2n$ ($f^{2}$ is not necessarily equal to I-{I}) is called a weak almost Hermitian manifold. If $\nabla^{g}J=0$, then such $(M^{2n},J,g)$ is called a weak Kähler manifold, see [9].
The $f^{2}$-torsion condition (3.3) in terms of a self-adjoint (1,1)-tensor $\widetilde{Q}=-f^{2}-{I}$ reads
@@ 4 Einstein connection of an NGT space
The following lemma is used in the proof of Theorem 4.3.
@@
Let $(M^{2n},f,g)$ be a weak almost Hermitian manifold, considered as a nonsymmetric pseudo-Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,fY)$. If an Einstein connection $\nabla$ on $M$ satisfies the $f^{2}$-torsion condition (3.3), then
$T(fY,fZ,({I}{-}f^{2})X)=$ $\,T(Y,Z,({I}{-}f^{2})f^{2}X){-}2(\nabla^{g}_{f^{2}X}\,F)(Y,fZ){+}2(\nabla^{g}_{f^{2}X}\,F)(fY,Z)$
$+dF(Y,Z,\,({I}-f^{2})f^{2}X)-dF(fY,fZ,\,({I}-f^{2})X),$ (4.1)
$T(fY,Z,\,({I}-f^{2})fX)=$ $\,-T(Y,fZ,\,({I}-f^{2})fX)-T(Y,Z,\,({I}-f^{4})X)$
$+dF(Y,Z,\,({I}-f^{4})X)+2\,(\nabla^{g}_{({I}-f^{2})X}\,F)(Y,Z)$
@@
The following theorem generalizes Theorem 2.8, it is valid for arbitrary $F$ satisfying (3.3), and for a non-degenerate tensor $P={I}-f^{2}$ we can use it to completely determine $T$.
@@
Let $\nabla$ be an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, satisfying the $f^{2}$-torsion condition (3.3), where $g(X,fY)=F(X,Y)$, and let the tensor $P:={I}-f^{2}$ have rank $2n$. If $fX\neq 0$, then the torsion $T$ of $\nabla$ is given by
@@
For $fX\neq 0$, it is carried out in a similar, but more computationally complex way than the proof of Theorem 2.8. Applying $Y\to fY$ and $Z\to fZ$ to (2.7), we have
@@
$\qquad-T(\widetilde{Q}Z,X,\widetilde{Q}Y)-T(X,\widetilde{Q}Y,\widetilde{Q}Z)-T(fY,\widetilde{Q}Z,fX)-T(\widetilde{Q}Y,fZ,fX).$ (4.6)
From (2.7) with $Y\to-f^{2}Y$, we get
@@
where $\delta_{4}(X,Y,Z)$ is given in Section 5. Using (2.8) and the $f^{2}$-torsion condition, we reduce $\delta_{4}(X,Y,Z)$ to $\delta_{5}(X,Y,Z)$, given in ection 5. Using (4.1) and (4.1), we find
@@
$T(Y,fZ,\widetilde{Q}f^{3}X)$ $+T(fY,Z,\widetilde{Q}f^{3}X)=dF(Y,Z,\widetilde{Q}f^{2}({I}+f^{2})X)+2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
@@
$+dF(Y,Z,(3\,\widetilde{Q}{+}2\,\widetilde{Q}^{2})f^{2}X){-}dF(Y,Z,\widetilde{Q}f^{2}({I}+f^{2})X){+}dF(Y,Z,\widetilde{Q}f^{2}X)$
@@
Take almost Hermitian manifolds $(M_{j},J_{j},g_{j})\ (j=1,\dots,m)$, where $J_{j}^{2}=-{I}_{\,j}$. Consider the Riemannian product $(M,g)=\bigl(M_{1}\times\cdots\times M_{m},\;g_{1}\oplus\cdots\oplus g_{m}\bigr)$, and define the (1,1)(1,1)–tensor $f=\sqrt{\lambda_{1}}\,J_{1}\oplus\cdots\oplus\sqrt{\lambda_{m}}\,J_{m}$. Then $(M,f,g)$ is a weak almost Hermitian manifold with $f^{2}=-\bigoplus_{j}\lambda_{j}\,{I}_{j}$, and on $TM_{j}$ (and for $X\in TM_{j}$) we have
@@
$\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,{I}_{j},\quad P={I}-f^{2}=\bigoplus\nolimits_{j}(1+\lambda_{j})\,{I}_{j},\quad P^{-1}=\bigoplus\nolimits_{j}\frac{1}{1+\lambda_{j}}\,{I}_{j}.$
@@
If an Einstein connection $\nabla$ satisfies the $f^{2}$-torsion condition (3.3), then $T(X,Y)=0$ for $X\in TM_{i},\,Y\in TM_{j}\ (i\neq j)$. Hence, the torsion tensor splits into the direct sum $T=T|_{TM_{1}}\oplus\cdots\oplus T|_{TM_{m}}$. Substituting $\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,{I}_{j}$ into (4.3), then using
@@
If each $(M_{j},J_{j},g_{j})$ is a Kähler manifold, then $\nabla^{g_{j}}F_{j}=0$ and $dF_{j}=0$. In this case, all torsion components vanish, and the Einstein connection reduces to the Levi-Civita connection. Using (2.5), we can further simplify the formula for $T|_{TM_{j}}$ as
$T|_{TM_{j}}(Y,Z,X)=$ $\ \frac{(\lambda-1)(5-3\lambda)}{4\lambda^{2}}\,(\nabla^{g_{j}}_{X}F_{j})(Y,Z)+\Bigl(\frac{1}{\lambda}+\frac{\lambda-1}{2\sqrt{\lambda}}\Bigr)\,(\nabla^{g_{j}}_{J_{j}X}F_{j})(Y,J_{j}Z)$
$-\Bigl(\frac{3\lambda^{2}-2\lambda+1}{4\lambda^{2}}+\frac{\lambda-1}{4\sqrt{\lambda}}\Bigr)\,\big\{(\nabla^{g_{j}}_{Z}F_{j})(X,Y)+(\nabla^{g_{j}}_{Y}F_{j})(Z,X)\big\}$
$+\Bigl(\frac{\lambda-1}{4\sqrt{\lambda}}{-}\frac{1}{2\lambda}{-}\frac{\lambda-1}{2\lambda^{2}}\Bigr)\,\big\{(\nabla^{g_{j}}_{J_{j}Y}F_{j})(J_{j}Z,X)+(\nabla^{g_{j}}_{J_{j}Z}F_{j})(X,J_{j}Y)\big\}.$
@@
Let $\nabla$ be an Einstein connection of a weak almost Hermitian manifold $(M^{2n},f,g>0)$, considered as a nonsymmetric Riemannian space $(M,G=g+F)$ with $F(X,Y)=g(X,fY)$. If the condition (3.3) is valid, then the torsion $T$ of $\nabla$ is given by (4.3).
@@
For a nonsymmetric Riemannian space $(M,G=g+F)$, the tensor $P={I}-f^{2}$ is positive definite, and therefore non-degenerate. Using (2.5), we can further simplify (4.3) to the form with $\nabla^{g}\,F$-terms.
@@
$T(fX,Y,Z)=-\tfrac{1}{3}\,dF(fX,Y,Z),\quad\nabla\,f^{2}=\nabla^{g}f^{2}=0.$